1. After having reviewed the lectures I have posted in Course Documerts perform the Consider Table (1) below. Using the data in this Table (3) calculate the value of r in the two formulas below the table. SHOW YOUR WORK AND OPERATIONS! TABLE 3) 14 12 10 2-103x x-bar and y-bar are MEANS of x and y n is the number of observations which is 8
r2 2Where, n is the mumber of observations in the sample (The data in Table 3). 2. Considering the values of e below: Calculate the value of d denoted in the formula below I cannot make typo corrections in the fomula, but instead Durbin-Watson Test Statistics
Answer #1

Answer:

1)

X

Y

(X-Mx)^2

(Y-My)^2

(X-Mx)(Y-My)

1

14

6.25

49

-17.5

1

12

6.25

25

-12.5

2

10

2.25

9

-4.5

3

8

0.25

1

-0.5

5

6

2.25

1

-1.5

5

4

2.25

9

-4.5

6

2

6.25

25

-12.5

5

0

2.25

49

-10.5

Mean

3.5

7

Sum

28

56

28

168

-64

Correlation coefficient:

ー64 ー-0.93314

X

Y

X^2

Y^2

xy

1

14

1

196

14

1

12

1

144

12

2

10

4

100

20

3

8

9

64

24

5

6

25

36

30

5

4

25

16

20

6

2

36

4

12

5

0

25

0

0

Mean

3.5

7

Sum

28

56

126

560

132

r=\frac{n\sum xy-\sum x\sum y\sum }{\sqrt{[n\sum x^2-(\sum x)^2][n\sum y^2-(\sum y)^2]}}=\frac{8*132-28*56}{\sqrt{[8*126-28^2][8*560-56^2]}}=-0.93314